# Edinburgh, EH9 3FB, UK https://andrewwango.github.io/perspective-equivariant-imaging

### v1

@@ Abstract
Ill-posed image reconstruction problems appear in many scenarios such as remote sensing, where obtaining high quality images is crucial for environmental monitoring, disaster management and urban planning. Deep learning has seen great success in overcoming the limitations of traditional methods. However, these inverse problems rarely come with ground truth data, highlighting the importance of unsupervised learning from partial and noisy measurements alone. We propose perspective-equivariant imaging (EI), a framework that leverages perspective variability in optical camera-based imaging systems, such as satellites or handheld cameras, to recover information lost in ill-posed optical camera imaging problems. This extends previous EI work to include a much richer non-linear class of group transforms and is shown to be an excellent prior for satellite and urban image data, where perspective-EI achieves state-of-the-art results in multispectral pansharpening, outperforming other unsupervised methods in the literature. Code at this URL.
@@
projective geometry inverse problems unsupervised learning image reconstruction remote sensing pansharpening
@@ 1 Introduction
Figure 1: Unsupervised pansharpening with perspective-equivariant imaging. Panchromatic (PAN) and low-res multispectral (MS) images are inputs. Reconstructions and average no-reference QNR (higher is better) [2] results for: upper 2 rows: commercially provided images from the bespoke classical HCS [48] method designed for WorldView-2 [41] which we treat as both a baseline and as oracle, linear upsampling baseline, oracle supervised training, self-supervised learning using Wald’s protocol [63]. Lower 2 rows: unsupervised loss functions from competitor methods [60, 18, 38, 39], shift-EI [13], and our method using our proposed unsupervised loss. All deep learning methods use the same neural network backbone for fair and balanced comparison of the loss functions; we are not comparing the different NN architectures from the literature. First 3 channels (RGB) are shown for visualisation. See Tab. 3(a) for more results.
@@
We propose perspective-equivariant imaging (EI), a framework that uses the natural belief that image sets are invariant to changes in perspective to solve these inverse problems without ground truth (GT). We formalise this by considering the group of non-linear projective transformations which goes beyond the simple linear transformations previously studied, by generalising to all possible transformations in a camera-based imaging system. Intuitively, we expect images taken from different perspective angles to belong to the same set of images, even if we don’t have access to the image set. We show that this information is enough to solve the inverse problem. Because the projective group is richer than the subgroups previously studied, we expect perspective-EI to learn a better mapping than existing EI methods. Note that we do not require image sequences as in structure-from-motion [47] nor multiple images of the same scene, just the assumption of image set invariance.
@@
The EI framework was proposed in [12] to learn to invert from measurements $\mathbf{y}$ alone. They showed that existing unsupervised methods that rely on measurement consistency (MC) alone (i.e. enforcing that $\mathbf{A}f(\mathbf{y})=\mathbf{y}$) fail to learn information about $\mathcal{X}$ outside the range of $\mathbf{A}^{\top}$. By postulating a mild prior that the unknown image set is invariant to a transformation group $G$, they construct an EI loss that allows the solver to "see" into the nullspace of the forward operator $\mathbf{A}$, and show necessary and sufficient conditions on $A$ and the group size $\lvert G\rvert$ in order for this to be possible [58]. The GT-free loss (see Fig. 2) is constructed as $\mathcal{L}_{\text{unsup}}(\theta)=\sum_{\mathbf{y}\in\{\mathbf{y}\}}\sum_{g\in G}\mathcal{L}_{\text{unsup}}(\theta;\mathbf{y},g)$ where:
@@
    $\left\{\mathbf{y}_{\text{MS}},\mathbf{y}_{\text{PAN}}\right\}\sim\mathcal{P}\left(\mathbf{A}_{\text{PS}}\mathbf{x}\right)=\left\{\mathcal{P}\left(\mathbf{A}_{\text{SR}}\mathbf{x}\right),\mathcal{P}\left(\mathbf{R}_{\text{PAN}}\mathbf{x}\right)\right\}$ (4)
    where $\mathbf{R}_{\text{PAN}}$ is the panchromatic (pan) channel’s spectral response function (SRF) and $\mathbf{A}_{\text{SR}}\mathbf{x}=(\mathbf{k}*\mathbf{x})\downarrow_{j}$ is the $j\times j$-factor downsampling operator with anti-aliasing kernel $\mathbf{k}$ (or modulation transfer function (MTF)). Structural information is lost in the downsampling and only one "band"’s worth of spectral information is kept in the pan band creating an ill-posed inverse problem. Satellite imaging is typically photon limited due to physical constraints on the aperture and therefore exhibits non-linear Poisson noise parameterised by the gain $\gamma$ (higher is noisier) [14].
The perspective-EI loss can also be used as an add-on to boost the performance of previous state-of-the-art unsupervised image reconstruction losses, is robust to noise and is network-agnostic and so can be used with the latest network architectures and training methods. Our method is easy to train (e.g. no adversarial training) and enjoys fast inference (i.e. no iteration).
@@ 1.1 Contributions
    We introduce a new invariance model for EI problems associated with the non-linear projective group structure of camera geometry, that is much richer than previously studied and goes beyond linear affine transformations;
@@ 2.0.1 Inverse problems in remote sensing
There are numerous ill-posed inverse problems actively studied in remote sensing image formation, as obtaining high quality satellite images is crucial, but physical constraints on satellites mean that there is a trade-off between the amount of captured spatial and spectral information. [42] Capturing images with fewer measurements $m$ also allows satellites to be more agile. This section attempts to unite the disjoint literature with the inverse problem framework.
Pansharpening (Eqs. 4 and 2(b)) is widely studied. Classical methods include component-substitution (CS), multiresolution analysis, and model-based variational optimisation approaches; see [42, 62] for surveys and benchmarks. Recently, several DL methods attempt to solve the inverse problem in an end-to-end manner [20]. However most of these are supervised models [74, 70, 68, 71, 69, 75] despite the unavoidable lack of available GT, and so must train their models with a self-supervised method by further degrading the LR input and treating it as the “ground truth” (the so-called Wald’s protocol [63]), assuming a scale invariance of the images, which is analogous to self-supervised methods in the superresolution and denoising literatures [55, 44]. However, they suffer from poor cross-scale generalisation when applied to the original input images [17] as we do not observe strong multi-scale self-similarity in satellite images. Furthermore, the noise model does not generalise from the high-res (HR) to the LR domain, reducing performance in noisy settings.
@@ 3 The projective transformation
Following [25, Sec. 8.4]. Two 2D images $\mathbf{x},\mathbf{x}^{\prime}$ taken of a 3D world from two cameras at arbitrary 3D orientations but the same coincident camera centre can be related via a projective transformation, written as a linear transformation in homogeneous coordinates. This can be visualised by moving the image plane in Fig. 4 but keeping the camera still.
@@
Our results also apply when two images are taken of a planar world from two cameras at different 3D locations/orientations [25, Sec. 13].
@@
Following [25, Sec. 8.4.2]. In Proposition 1, all of the non-linear projective transformations ("homographies") can be written as a linear transformation $\mathbf{\bar{x}}^{\prime}=\mathbf{T}_{g}\mathbf{\bar{x}},\mathbf{T}_{g}\in\mathbb{R}^{3\times 3}$ drawn from the group $g\in G$, where $\bar{\mathbf{x}}=(u,v,1)^{\top},\bar{\mathbf{x}}^{\prime}=(u^{\prime},v^{\prime},1)^{\top}$ are the image homogeneous coordinates. Note that while the homography is linear in homogeneous coordinates, it is generally a highly non-linear transform [25, Sec. 2.3]. This can be decomposed as
@@
where $\mathbf{K},\mathbf{K}^{\prime}$ are the intrinsics of the cameras before and after transformation (see Fig. 4 for definitions). Note that for most usual cameras $m_{x}=m_{y}=1,s=0$ [25, Sec. 6.2.4], but we will relax this when describing affine cameras. $\mathbf{R}^{\prime}=\mathbf{R}_{z}(\theta_{z})\mathbf{R}_{y}(\theta_{y})\mathbf{R}_{x}(\theta_{x})$ is the relative 3D rotation matrix of the image plane about the camera centre $C$, parametrised by the Euler angles of rotation $\theta_{x},\theta_{y},\theta_{z}$.
@@ 3.1 Group structure
Homographies defined above are the actions of the projective linear group $\text{PGL}(3)$ [40] which gives rise to a hierarchical subgroup structure [25, Sec. 2.4], summarised in Tab. 1. We apply the homography supergroup for perspective-EI, which has substantially more symmetries, as originally illustrated in Klein geometry theory [32]. Intuitively, this means that we can expect to learn more information from the nullspace, and further constrain the structure of $f_{\theta}$ therefore also making training more data-efficient [13, 58]. In a geometrical interpretation, each subgroup’s transformations can be recovered by fixing certain parameters in the homography and compound transforms can be made by intersecting these parameterisations; see SM for derivations. We have a hierarchy of transformations where we can choose an appropriate subgroup $G$ depending on the application, as long as $A$ is itself not $G$-equivariant [13]. The homography has more general applicability since there will be fewer $A$s that are $G$-equivariant. For our experiments we will compare the various parametrisations from Tab. 1 - although the full similarity and affine transformations do not contain perspective shifts, they are also larger than previously studied simpler transforms.
@@
Transform
Ref
Fix all parameters except…
shift
[13]
principal point $u^{(0)\prime},v^{(0)\prime}$
rotation
[13]
$\theta_{z}^{\prime}$
scale
[53]
focal length $f^{\prime}$
similarity (shift+rot+scale)
ours
$f^{\prime},u^{(0)\prime},v^{(0)\prime},\theta_{z}^{\prime}$
affine (similarity+skew)
ours
as above except $m_{x}^{\prime},m_{y}^{\prime},s^{\prime}$
pan and tilt
ours
$\theta_{x}^{\prime},\theta_{y}^{\prime}$
perspective (similarity+pan+tilt)
ours
as similarity except $\theta_{x}^{\prime},\theta_{y}^{\prime}$.
@@ 3.1.1 Pan and tilt transformation
[…]
@@ 3.1.2 Proposed method
Our perspective-EI model is shown in Fig. 2, which is agnostic to choice of NN $f_{\theta}$ e.g. feed-forward nets, CNNs, attention-based models, unrolled models etc. The unsupervised loss consists of an MC loss and perspective equivariance loss Eq. 2, where the group transformation is given in Tab. 1. Note the MC loss can be any loss depending on the specific inverse problem, e.g. the MSE in Eq. 2, L1 [73], SURE in noisy data (Eq. 3), or specific MC losses for multispectral imaging; see Sec. 4.2.
@@
Method PSNR SSIM MC 7.14 0.076 Oracle supervised 24.97 0.838 Shift EI 26.47 0.865 Rotation EI 26.29 0.861 Pan+tilt EI (ours) 27.52 0.887 Similarity EI (ours) 27.63 0.89 Perspective-EI (ours) 27.67 0.889
(b) Average test PSNR and SSIM.
@@ 4.1 Image inpainting
The backbone NN is a U-Net [51], used for EI in [13], but the choice of NN is independent of the comparison of our loss functions. We compare the transformations from Tab. 1, and compare to baselines: pure MC (i.e. $\mathcal{L}_{\text{unsup}}=\mathcal{L}_{\text{MSE}}$), which cannot recover information from the nullspace of the operator, and "oracle" supervised where $\mathcal{L}_{\text{sup}}(\theta;\mathbf{x},\mathbf{y})=\lVert f_{\theta}(\mathbf{y})-\mathbf{x}\rVert_{2}^{2}$ which has access to the paired GT images $\mathbf{x}$. See SM for additional experimental details.
Quantitative and qualitative results are shown in Fig. 6. All EI methods are able to learn into the nullspace and achieve excellent results compared to MC with all of the transforms. However we observe significantly better results when using our proposed pan+tilt non-linear transform, compared to the linear shifts and rotations studied in [13]. Furthermore, we outperform the above when using compound transforms from Tab. 1 such as similarity (composed of shift, rotation and scale) and perspective - while simpler transforms are able to uncover the nullspace information, increasing the structure in the model allows us to train better with less data. Note that supervised training performs badly on the test set due to a high generalisation gap; see SM for more results. We also observe better performance when restricting to a 10% range of transformations rather than the full range; for the perspective transformation, this avoids singularities and interpolation distortions at infinity. See ablation study in SM for comparisons.
@@ 4.2 Multispectral satellite image pansharpening
We train and test on a random split of 957 vs 107 $256\times 256\times 4$ RGB+NIR LR-MS and $1024\times 1024$ pan tiles from a subset of the SpaceNet-4 [66] dataset where all images are labelled at $42\degree$ off-nadir. $\mathbf{y}_{\text{MS}}$ tiles were simulated by downsampling and noising the WorldView-2 [41] HR-MS tiles in order to simulate a known Gaussian MTF, and avoid registration errors. Note these HR-MS tiles were commercially provided by WorldView-2, pansharpened with a bespoke HCS [48] method designed for WorldView-2. While they might not represent ideal GT spectral and structural fidelity, we use this as a classical baseline to which we compare the DL methods on quality metrics. The $\mathbf{y}_{\text{pan}}$ tiles are used as provided i.e. from an unknown SRF. The scenes are highly challenging as they cover a diverse range of urban and rural areas.
@@
Paper
Loss function
Notes
GDD [60]
$\lVert\ldots\rVert_{2}^{2}+\lVert\ldots\rVert_{\text{TV}}$
i.e. our pure MC baseline
PanGan [39]
$\lVert\ldots\rVert_{2}^{2}+\lVert\nabla\ldots\rVert_{2}^{2}+\mathcal{L}_{\text{adv}}(\mathbf{y},\hat{\mathbf{x}})$
where $\mathcal{L}_{\text{adv}}$ is an adversarial consistency
Z-PNN [18]
$\lVert\ldots\rVert_{1}+\text{corr}(\mathbf{y}_{\text{pan}},f_{\theta}(\mathbf{y}))$
where corr is a spatial correlation
SSQ [38]
$\mathfrak{L}(\mathbf{k}*\mathbf{\hat{x}},\mathbf{y}_{\text{MS}\uparrow})+\mathfrak{L}(\mathbf{R}_{\text{pan}}\mathbf{\hat{x}},\mathbf{y}_{\text{pan}})+1-\text{QNR}(\mathbf{\hat{x}})$
where $\mathfrak{L}$ is a combination of MSE and SSIM and $\mathbf{y}_{\text{MS}\uparrow}$ is an upsampling.
[16] offers a recent benchmark and shows that the losses in [60, 38, 39, 18] are all forms of MC losses. While they compare the full models directly, we only compare their loss functions and use the same NN backbone for all models for a fair comparison, as the choice of unsupervised loss is independent of the choice of NN (i.e. our experiments are not comparable with the original papers where different NNs were used). We use PanNet [70] which is lightweight but competitive in benchmarks [20, 16]. We do not compare with [16] as it is essentially [18] but with a more complex spectral loss which could be easily appended to any method. [60, 38, 39] all also differ in their estimation of $\mathbf{R}_{\text{pan}}$: [60] jointly learns it, [38] estimates at LR using linear regression before training, and [39] fixes it as an average. Here, we use the same flat $\mathbf{R}_{\text{pan}}$ for a fair comparison across methods, as we find this performs best; see SM for comparisons.
We also compare against traditional self-supervised learning i.e. Wald’s protocol [63] and supervised learning using the commercially provided WorldView-2 bespoke HCS pansharpened images as oracle. We do not compare against DIP [61] as it has already been shown to not be competitive in [60]. We report the popular quality with no reference (QNR) metric [2] to measure spectral and structural pansharpening quality; see SM for details. For the noisy case we do not report QNR as it cannot measure denoising capabilities. Although there is no real HR-MS GT (recall that these are commercially provided by WorldView-2), we also report PSNR and ERGAS [62] as a "sense-check" compared to the commercially-provided images pansharpened with the bespoke CS method, HCS. Finally, for the noisy case we also report a baseline using Brovey [62], a classical CS method like HCS, widely used in practice [42].
@@
Table 3: Quantitative results for pansharpening on the SpaceNet-4 test set (qualitative in Figs. 1 and 7) with best unsupervised method in bold. Average PSNR and ERGAS are compared to the commercially-provided reference images pansharpened with the bespoke HCS. QNR is no-reference. For descriptions of competitors see Sec. 4.2.
@@
Method QNR $\uparrow$ PSNR $\uparrow$ ERGAS $\downarrow$ Baseline (bespoke HCS) 0.87 $\infty$ 0 Baseline (linear) 0.142 7.01 24.83 (Oracle supervised) (0.857) (25.61) (3.71) Self-sup (Wald) 0.796 20.03 6.95 Competitor: GDD 0.723 22.48 4.86 Competitor: SSQ 0.627 21.28 5.46 Competitor: Z-PNN 0.798 22.79 4.89 Competitor: PanGan 0.836 20.56 6.63 Shift EI 0.866 25.02 3.71 Affine EI (ours) 0.88 25.29 3.71 Pan+tilt EI (ours) 0.872 25.4 3.57 Perspective-EI (ours) 0.892 25.16 3.82

### v2

# %****␣main.tex␣Line␣75␣****%TODO␣FINAL:␣Replace␣with␣your␣institution␣list.https://andrewwango.github.io/perspective-equivariant-imaging

@@ Abstract
Ill-posed image reconstruction problems appear in many scenarios such as remote sensing, where obtaining high quality images is crucial for environmental monitoring, disaster management and urban planning. Deep learning has seen great success in overcoming the limitations of traditional methods. However, these inverse problems rarely come with ground truth data, highlighting the importance of unsupervised learning from partial and noisy measurements alone. We propose perspective-equivariant imaging (EI), a framework that leverages classical projective camera geometry in optical imaging systems, such as satellites or handheld cameras, to recover information lost in ill-posed camera imaging problems. We show that our much richer non-linear class of group transforms, derived from camera geometry, generalises previous EI work and is an excellent prior for satellite and urban image data. Perspective-EI achieves state-of-the-art results in multispectral pansharpening, outperforming other unsupervised methods in the literature. Code at this URL.
@@
projective geometry inverse problems unsupervised learning image reconstruction remote sensing
@@ 1 Introduction
Figure 1: Unsupervised pansharpening with perspective-equivariant imaging. Panchromatic (PAN) and low-res multispectral (MS) images are inputs. Reconstructions and average no-reference QNR (higher is better) [2] results for: upper 2 rows: commercially provided images from the bespoke classical Hyperspherical Color Space (HCS) [48] method designed for WorldView-2 [41] which we treat as both a baseline and as oracle, linear upsampling baseline, oracle supervised training, self-supervised learning using Wald’s protocol [63]. Lower 2 rows: unsupervised loss functions from competitor methods [60, 18, 38, 39], shift-EI [13], and our method using our proposed unsupervised loss. All deep learning methods use the same neural network backbone for fair and balanced comparison of the loss functions; we are not comparing the different NN architectures from the literature. First 3 channels (RGB) are shown for visualisation. See Tab. 3(a) for more results.
@@
We propose perspective-equivariant imaging (EI), a framework that uses the natural belief that image sets are invariant to changes in perspective to solve these inverse problems without ground truth (GT). We formalise this by considering the group of non-linear projective transformations (homographies) derived from traditional camera geometry. Intuitively, we expect images taken from different perspective angles to belong to the same set of images, even if we don’t have access to the image set. We show that this information is enough to solve the inverse problem. This goes beyond the simple linear transformation subgroups previously studied, by generalising to all possible transformations in a camera-based imaging system, and we expect perspective-EI to learn a better mapping than existing EI methods. Note that we do not require image sequences as in structure-from-motion [47] nor multiple images of the same scene, just the assumption of image set invariance.
@@
The EI framework was proposed in [12] to learn to invert from measurements $\mathbf{y}$ alone. They showed that existing unsupervised methods that rely on measurement consistency (MC) alone (i.e. enforcing that $\mathbf{A}f(\mathbf{y})=\mathbf{y}$) fail to learn information about $\mathcal{X}$ outside the range of $\mathbf{A}^{\top}$. By postulating a mild prior that the unknown image set is invariant to a transformation group $G$, they construct an EI loss that allows the solver to "see" into the nullspace $\mathbf{x}:\mathbf{Ax}=\mathbf{0}$ of the forward operator $\mathbf{A}$, and show necessary and sufficient conditions on $\mathbf{A}$ and the group size $\lvert G\rvert$ in order for this to be possible [58]. The GT-free loss (see Fig. 2) is constructed as $\mathcal{L}_{\text{unsup}}(\theta)=\sum_{\mathbf{y}\in\{\mathbf{y}\}}\sum_{g\in G}\mathcal{L}_{\text{unsup}}(\theta;\mathbf{y},g)$ where:
@@
    $\left\{\mathbf{y}_{\text{MS}},\mathbf{y}_{\text{PAN}}\right\}\sim\mathcal{P}\left(\mathbf{A}\mathbf{x}\right)=\left\{\mathcal{P}\left(\mathbf{A}_{\text{SR}}\mathbf{x}\right),\mathcal{P}\left(\mathbf{R}_{\text{PAN}}\mathbf{x}\right)\right\}$ (4)
    where MS = multispectral, PAN = panchromatic, $\mathbf{R}_{\text{PAN}}$ is the pan channel’s spectral response function (SRF) and $\mathbf{A}_{\text{SR}}\mathbf{x}=(\mathbf{k}*\mathbf{x})\downarrow_{j}$ is the $j\times j$-factor downsampling operator with anti-aliasing kernel $\mathbf{k}$ (or modulation transfer function (MTF)). Structural information is lost in the downsampling and only one "band"’s worth of spectral information is kept in the pan band creating an ill-posed inverse problem. Satellite imaging is typically photon limited due to physical constraints on the aperture and therefore exhibits non-linear Poisson noise parameterised by the gain $\gamma$ (higher is noisier) [14].
The perspective-EI loss is network-agnostic so can be used alongside the latest network architectures and training methods; it can also be used as an add-on to boost the performance of previous state-of-the-art unsupervised image reconstruction losses. Our method is robust to noise and easy to train (e.g. no adversarial training) and enjoys fast inference (i.e. no iteration).
@@ 1.1 Contributions
    We introduce a new, larger group invariance model for equivariant imaging problems using the non-linear projective group structure of camera geometry, enabling new unsupervised learning for a wider range of imaging tasks;
@@ 2.0.1 Inverse problems in remote sensing
There are numerous ill-posed inverse problems actively studied in remote sensing image formation, as obtaining high quality satellite images is crucial, but physical constraints on satellites mean that there is a trade-off between the amount of captured spatial and spectral information [42]. Capturing images with fewer measurements $m$ also allows satellites to be more agile. This section attempts to unite the disjoint literature with the inverse problem framework.
Pansharpening (Eqs. 4 and 3(b)) is widely studied. Classical methods include component-substitution (CS), multiresolution analysis, and model-based variational optimisation approaches; see [42, 62] for surveys and benchmarks. Recently, several DL methods attempt to solve the inverse problem in an end-to-end manner [20]. However most of these are supervised models [74, 70, 68, 71, 69, 75] despite the unavoidable lack of available GT, and so must train their models with a self-supervised method by further degrading the LR input and treating it as the “ground truth” (the so-called Wald’s protocol [63]), assuming a scale invariance of the images, which is analogous to self-supervised methods in the superresolution and denoising literatures [55, 44]. However, they suffer from poor cross-scale generalisation when applied to the original input images [17] as we do not observe strong multi-scale self-similarity in satellite images. Furthermore, the noise model does not generalise from the high-res (HR) to the LR domain, reducing performance in noisy settings.
@@ 3 The projective transformation
Following [26, Sec. 8.4]. Two 2D images $\mathbf{x},\mathbf{x}^{\prime}$ taken of a 3D world from two cameras at arbitrary 3D orientations but the same coincident camera centre can be related via a projective transformation, written as a linear transformation in homogeneous coordinates. This can be visualised by moving the image plane in Fig. 4 but keeping the camera still.
@@
Our results also apply when two images are taken of a planar world from two cameras at different 3D locations/orientations [26, Sec. 13].
@@
Following [26, Sec. 8.4.2]. In Proposition 1, all of the non-linear projective transformations ("homographies") can be written as a linear transformation $\mathbf{\bar{x}}^{\prime}=\mathbf{T}_{g}\mathbf{\bar{x}},\mathbf{T}_{g}\in\mathbb{R}^{3\times 3}$ drawn from the group $g\in G$, where $\bar{\mathbf{x}}=(u,v,1)^{\top},\bar{\mathbf{x}}^{\prime}=(u^{\prime},v^{\prime},1)^{\top}$ are the image homogeneous coordinates. Note that while the homography is linear in homogeneous coordinates, it is generally a highly non-linear transform [26, Sec. 2.3]. This can be decomposed as
@@
where $\mathbf{K},\mathbf{K}^{\prime}$ are the intrinsics of the cameras before and after transformation (see Fig. 4 for definitions). Note that for most usual cameras $m_{x}=m_{y}=1,s=0$ [26, Sec. 6.2.4], but we will relax this when describing affine cameras. $\mathbf{R}^{\prime}=\mathbf{R}_{z}(\theta_{z})\mathbf{R}_{y}(\theta_{y})\mathbf{R}_{x}(\theta_{x})$ is the relative 3D rotation matrix of the image plane about the camera centre $C$, parametrised by the Euler angles of rotation $\theta_{x},\theta_{y},\theta_{z}$.
@@ 3.1 Group structure
Homographies defined above are the actions of the projective linear group $\text{PGL}(3)$ [40] which gives rise to a hierarchical subgroup structure [26, Sec. 2.4], summarised in Tab. 1. We apply the homography supergroup for perspective-EI, which has substantially more symmetries, as originally illustrated in Klein geometry theory [32]. Intuitively, this means that we can expect to learn more information from the nullspace, and further constrain the structure of $f_{\theta}$ therefore also making training more data-efficient [13, 58]. In a geometrical interpretation, each subgroup’s transformations can be recovered by fixing certain parameters in the homography and compound transforms can be made by intersecting these parameterisations; see SM for derivations. We create a hierarchy of transformations where we can choose an appropriate subgroup $G$ depending on the application, as long as $A$ is itself not $G$-equivariant [13]. The homography has more general applicability since there will be fewer $A$s that are $G$-equivariant. For our experiments we will compare the various parametrisations from Tab. 1 - although the full similarity and affine transformations do not contain perspective shifts, they are also larger than previously studied simpler transforms.
@@
Transform Ref Fix all parameters except…
shift [13] principal point $u^{(0)\prime},v^{(0)\prime}$
rotation [13] $\theta_{z}^{\prime}$
scale [53] focal length $f^{\prime}$
similarity (shift+rot+scale) ours $f^{\prime},u^{(0)\prime},v^{(0)\prime},\theta_{z}^{\prime}$
affine (similarity+skew) ours as above except $m_{x}^{\prime},m_{y}^{\prime},s^{\prime}$
pan and tilt ours $\theta_{x}^{\prime},\theta_{y}^{\prime}$
perspective (similarity+pan+tilt) ours as similarity except $\theta_{x}^{\prime},\theta_{y}^{\prime}$.
@@ 3.1.1 Pan and tilt transformation
form a subgroup $G=(\mathbf{\Pi},\cdot)$ of the group of general homographies, defined on the projective space $\mathbb{P}^{2}$. Furthermore, $\mathbf{\Pi}$ is isomorphic to the 3D rotation group on $\mathbb{R}^{3}$, the special orthogonal group $\text{SO}(3)$. Proofs: see SM.
@@ 3.1.2 Proposed method
Our perspective-EI model is shown in Fig. 2, which is agnostic to choice of NN $f_{\theta}$ e.g. feed-forward nets, CNNs, attention-based models, unrolled models etc. The unsupervised loss consists of an MC loss and perspective equivariance loss Eq. 2, where the group transformation is Eq. 5 constructed from projective goemetry. Note the MC loss can be any loss depending on the specific inverse problem, e.g. the MSE in Eq. 2, L1 [73], SURE in noisy data (Eq. 3), or specific MC losses for multispectral imaging; see Sec. 4.2.
@@
(b) Average test PSNR and SSIM. Method PSNR SSIM MC 7.14 0.076 Oracle supervised 24.97 0.838 Shift EI 26.47 0.865 Rotation EI 26.29 0.861 Pan+tilt EI (ours) 27.52 0.887 Similarity EI (ours) 27.63 0.89 Perspective-EI (ours) 27.67 0.889
@@ 4.1 Image inpainting
The backbone NN is a U-Net [51], used for EI in [13], but the choice of NN is independent of the comparison of our loss functions. We compare the transformations from Tab. 1, and compare to baselines: pure MC (i.e. $\mathcal{L}_{\text{unsup}}=\mathcal{L}_{\text{MSE}}$), which cannot recover any more information than identity from the nullspace, and "oracle" supervised where $\mathcal{L}_{\text{sup}}(\theta;\mathbf{x},\mathbf{y})=\lVert f_{\theta}(\mathbf{y})-\mathbf{x}\rVert_{2}^{2}$ which has access to the paired GT images $\mathbf{x}$. See SM for additional experimental details.
Quantitative and qualitative results are shown in Fig. 6. All EI methods are able to learn into the nullspace and achieve excellent results compared to MC with all of the transforms. However we observe significantly better results when using our proposed pan+tilt non-linear transform, compared to the linear shifts and rotations studied in [13]. Furthermore, we outperform the above when using compound transforms from Tab. 1 such as similarity (composed of shift, rotation and scale) and perspective - while simpler transforms are able to uncover the nullspace information, increasing the structure in the model allows us to train a stronger model. Note that supervised training performs badly on the test set due to a high generalisation gap; see SM for more results. We also observe better performance when restricting to a 10% range of transformations rather than the full range; for the perspective transformation, this avoids singularities and interpolation distortions at infinity. See ablation study in SM for comparisons.
@@ 4.2 Multispectral satellite image pansharpening
We train and test on a random split of 957 vs 107 $256\times 256\times 4$ RGB+NIR LR-MS and $1024\times 1024$ pan tiles from a subset of the SpaceNet-4 [66] dataset where all images are labelled at $42\degree$ off-nadir. $\mathbf{y}_{\text{MS}}$ tiles were simulated by downsampling and noising the WorldView-2 [41] HR-MS tiles in order to simulate a known Gaussian MTF, and avoid registration errors. Note these HR-MS tiles were commercially provided by WorldView-2, pansharpened with a bespoke Hyperspherical Color Space (HCS) [48] method designed for WorldView-2. While they might not represent ideal GT spectral and structural fidelity, we use this as a classical baseline to which we compare the DL methods on quality metrics. The $\mathbf{y}_{\text{pan}}$ tiles are used as provided i.e. from an unknown SRF. The scenes are highly challenging as they cover a diverse range of urban and rural areas.
@@
Paper Loss function Notes
GDD [60] $\lVert\ldots\rVert_{2}^{2}+\lVert\ldots\rVert_{\text{TV}}$ i.e. our pure MC baseline
PanGan [39] $\lVert\ldots\rVert_{2}^{2}+\lVert\nabla\ldots\rVert_{2}^{2}+\mathcal{L}_{\text{adv}}(\mathbf{y},\hat{\mathbf{x}})$ where $\mathcal{L}_{\text{adv}}$ is an adversarial consistency
Z-PNN [18] $\lVert\ldots\rVert_{1}+\text{corr}(\mathbf{y}_{\text{pan}},f_{\theta}(\mathbf{y}))$ where corr is a spatial correlation
SSQ [38] $\mathfrak{L}(\mathbf{k}*\mathbf{\hat{x}},\mathbf{y}_{\text{MS}\uparrow})+\mathfrak{L}(\mathbf{R}_{\text{pan}}\mathbf{\hat{x}},\mathbf{y}_{\text{pan}})+1-\text{QNR}(\mathbf{\hat{x}})$ where $\mathfrak{L}$ is a combination of MSE and SSIM and $\mathbf{y}_{\text{MS}\uparrow}$ is an upsampling.
[16] offers a recent benchmark and shows that the losses in [60, 38, 39, 18] are all forms of MC losses. While they compare the full models directly, we only compare their loss functions and use the same NN backbone for all models for a fair comparison, as the choice of unsupervised loss is independent of the choice of NN (i.e. our experiments are not comparable with the original papers where different NNs were used). We use PanNet from the original paper [70] which is lightweight but competitive in benchmarks [20, 16]. We do not compare with [16] as it is essentially [18] but with a more complex spectral loss which could be easily appended to any method. [60, 38, 39] all also differ in their estimation of $\mathbf{R}_{\text{pan}}$: [60] jointly learns it, [38] estimates at LR using linear regression before training, and [39] fixes it as an average. Here, we use a flat $\mathbf{R}_{\text{pan}}$ for a fair comparison across methods, as we find this performs best; see SM for comparisons.
We also compare against traditional self-supervised learning i.e. Wald’s protocol [63] and supervised learning using the commercially provided WorldView-2 bespoke HCS pansharpened images as oracle. We do not compare against DIP [61] as it has already been shown to not be competitive in [60]. We report the popular quality with no reference (QNR) metric [2] to measure both spectral and structural distortion for pansharpening quality; see SM for details. For the noisy case we do not report QNR as it cannot measure denoising capabilities. Although there is no real HR-MS GT, we also report PSNR, ERGAS [62] and Spectral Angle Mapper (SAM) [42] as a "sense-check" compared to the images commercially-provided by WorldView-2 pansharpened with the bespoke CS method, HCS. Finally, for the noisy case we also report a baseline using Brovey [62], a classical CS method like HCS, widely used in practice [42].
@@
Table 3: Quantitative results for pansharpening on the SpaceNet-4 test set (qualitative in Figs. 1 and 7) with best unsupervised method in bold. Mean PSNR, ERGAS [62], SAM [42] are compared to commercially-provided reference images pansharpened with the bespoke HCS. QNR is no-reference. For descriptions of competitors see Sec. 4.2.
@@
Method QNR $\uparrow$ PSNR $\uparrow$ ERGAS $\downarrow$ SAM $\downarrow$ Baseline (bespoke HCS) 0.87 $\infty$ 0 0 Baseline (linear) 0.142 7.01 24.83 1.017 (Oracle supervised) 0.857 (25.61) (3.71) (0.080) Self-sup (Wald) 0.796 20.03 6.95 0.146 Competitor: GDD 0.723 22.48 4.86 0.121 Competitor: SSQ 0.627 21.28 5.46 0.141 Competitor: Z-PNN 0.798 22.79 4.89 0.117 Competitor: PanGan 0.836 20.56 6.63 0.325 Shift EI 0.866 25.02 3.71 0.086 Affine EI (ours) 0.88 25.29 3.71 0.084 Pan+tilt EI (ours) 0.872 25.4 3.57 0.082 Perspective-EI (ours) 0.892 25.16 3.82 0.085